A piggy- bank contains 50 coins, all either 2 cents or 5 cents. The total value of coins is $1.87. How many 2 cents coins are there?
Answer by april4869(19)

Let x represent the number of 2-cent coins.
Therefore, 50 - x = number of 5-cent coins.
Value of bank in cents is 2x + 5(50 - x) = 187 (notice every value is in cents) solve for x.
-3x + 250 = 187
-3x = 187 - 250 = -63
x = -63/-3 = 21
There are 21 2-cent coins and 29 5-cent coins.